# Y-AR: A Mixed Reality CAD Tool for 3D Wire Bending

### v1

@@ 8.6. Results
User Experience. All ratings in this evaluation are based on a 10-point scale and result is shown in the Figure 15. Out of the 12 participants, 9 indicated they would like to use Y-AR frequently, rating it more than 7. Ten participants felt confident using Y-AR, also rating it more than 7, and 11 believed most people would quickly learn to use Y-AR, with ratings above 7. Additionally, 10 participants enjoyed using hand gesture modeling, rating this aspect above 7, and 10 enjoyed the design process overall, with ratings above 7; 2 participants rated it at 6. The AR environment was found useful for obtaining real-world dimensions by all participants. Eight participants appreciated the gestural icons, rating them more than 7. Only three participants found Y-AR complex, rating this aspect above 3, and only two participants indicated they needed to learn many things before starting, rating this aspect above 7. Lastly, 11 participants considered the environment of the room in their design, with ratings above 7.
Technical Performance. As shown in Figure 16, the accuracy of hand gesture recognition was rated positively by 6 participants, with scores above 7, and 3 participants rated it at 6. The mechanical connector design met the expectations of 10 participants, who rated it above 7, while 2 participants rated it at 6. Despite only 4 participants having a clear strategy in mind when designing the connector, all were able to design a connector successfully using Y-AR, with 4 rating their strategy above 7 and 3 at 6. Eleven participants reported a good sense of size, with ratings above 7, and similarly, 11 participants felt confident in the spatial positioning while using Y-AR. Furthermore, all participants believed that the connectors, once manufactured, would look as expected, rating this aspect above 7, and 11 participants felt the connectors would function as intended, also rating above 7. Additionally, all participants acknowledged that Y-AR helped them design a functional object.
Areas for Improvement. As shown in Figure 17, the results also highlighted some areas for improvement. Three participants felt they would need technical support to use Y-AR effectively, rating this need less than 3, and 2 participants rated it at 4, suggesting that Y-AR may require a more intuitive on-boarding process. Additionally, four participants reported experiencing arm fatigue, with scores indicating no arm fatigue below 4.
ubsectionDiscussion
Ease of Use and Intuitiveness.
Participants found Y-AR straightforward and easy to use, highlighting its simplicity and intuitive design. For example, P12 mentioned, ”It was easy to make… So it’s really useful for me to use pre-made wire same.” P11 appreciated Y-AR’s straightforwardness: ”I think using the gesture could easily reflect the actual size and is very straightforward without the hassles of using measuring rulers to measure the wires and then put them into the CAD model.”
P10 compared Y-AR favorably to other drawing tools, stating, ”Unlike some other drawing tools, where I have to select different pens, thickness, and colors, I can just draw and get what I want right away.” P8 initially felt overwhelmed due to unfamiliarity with the XR environment but noted, ”After some kind of introduction, I was getting familiar with that. I was pretty comfortable during the experiment.”
P7 found the freehand drawing feature intriguing, saying it was ”very interesting to be able to draw freely in the space around me.” However, P6 highlighted a potential issue, expressing that while Y-AR is simple, ”from my perspective, it doesn’t seem that user-friendly… especially with the headset’s operation logic and gesture recognition,” suggesting a need for better support in understanding XR operation systems.
P3 enjoyed Y-AR’s features, remarking, ”I really enjoy the function of the tool as you can draw anything you want in the VR world, which is pretty interesting, amazing.” They also found it helpful for connecting different parts and having dimensions available, which they described as ”simple and helpful.” Finally, P1 succinctly captured Y-AR’s appeal, stating, ”The tool is very intuitive, allowing users to quickly understand what they should do.”
Real-Time Feedback and Visualization.
Participants appreciated Y-AR’s ability to provide immediate feedback on designs and see changes in the structure in real-time, finding this feature valuable for making quick adjustments and ensuring their designs met expectations. For instance, P6 noted, ”This feature is particularly helpful for designing objects that have practical use in real life. It provides a significant visual aid during the design process.”
P6 highlighted the benefits of using XR devices for design, contrasting it with traditional tools like Blender: ”Using tools like Blender, I found it difficult to visualize 3D space. However, with VR, seeing a 3D image projected into the space provides significant help, especially in practical design applications. For instance, even if I know the dimensions of a table, seeing it visually in the space makes a big difference.”
Potential for Practical Application.
Participants saw significant potential for Y-AR in various practical applications. They envisioned using it for creating customized holders or organizers, particularly in home and kitchen settings. P11 noted, ”I think for wires like charging cable… we definitely need organizers to make everything in order, so I think that is a good use case scenario.” The tool was also recognized for its potential in artistic and creative applications. P9 discussed this, mentioning, ”I can definitely imagine this to be for not their practical but more artistic purpose… just hanging on your wall simply.”
Participants highlighted several applications, such as phone holders and organizers for items. P8 shared, ”For normal people, they would like to use a phone holder. It’s a very common use case. And also, something like attaching a bottle to the table or similar applications.” P8 also suggested a feature enhancement: ”Probably we can add something like if you have a phone and it’s set at a certain angle, it could automatically generate a structure underneath to support it.”
Feature Requests
Participants provided several suggestions for enhancements and new features in Y-AR to improve the user experience and functionality in the future. One key recommendation was the inclusion of more templates, which would assist users in creating more complex or functional designs without needing detailed technical knowledge. As one participant expressed, ”Having more predefined structures would be really helpful” (P12). Another participant highlighted the potential benefit of a ”measurement feature” similar to those in CAD tools, which would allow users to easily see dimensions and scale objects (P7).
9. CONCLUSION AND FUTURE WORK
@@
We presented Y-AR, a CAD tool for 3D wire bending. This fabrication-aware environment allows designers to create structures with good wire bending in mind. We see our guidelines for design for wire bending as contributions independent of the specific implementation of the Y-AR environment. The core insight to leverage the springiness of the material to overcome modeling inaccuracies allows Y-AR to rely on much less precise input techniques such as mid-air gestures in XR. This XR environment, in turn, contributes to better design by guiding users to make objects that fit in with their context of use, we imagine this central idea will allow developers to build other fabrication-aware CAD tools for other fabrication techniques than wire bending. Our user evaluation indicates that Y-AR indeed allows users to create functional structures with wire bending that would have been hard to accomplish with existing CAM tools for wire bending.
@@
Y-AR will be available as an open source tool for other researchers to extend. It would be interesting to extend Y-AR with new types of tools, or consider facilitating design of structures that are complemented by other fabrication techniques than wire bending alone. Related work has shown that wire bending is well equipped to provide structure for other fabrication processes. We are also interested in running more longer-term deployment to see how users benefit from Y-AR over an extended amount of time. And as stated before, the framework of Y-AR combining XR and springiness could form a backbone for other future fabrication-aware design environments.
@@ Acknowledgements.
We thank the XR Collaboratory at Cornell Tech for their generous sponsorship of this research in the form of the inaugural XR Prototyping Grant.References
@@
- Hsiao et al. (2018) Kai-Wen Hsiao, Jia-Bin Huang, and Hung-Kuo Chu. 2018. Multi-view wire art. ACM Trans. Graph. 37, 6, Article 242 (dec 2018), 11 pages. https://doi.org/10.1145/3272127.3275070
- Huang et al. (2023) Huiguang Huang, Dong-Yi Wu, Hanh Le, Po-Chih Chen, Shih-Syun Lin, and Tong-Yee Lee. 2023. Learning to Generate Wire Sculpture Art from 3D Models. In SIGGRAPH Asia 2023 Posters (Sydney, NSW, Australia) (SA ’23). Association for Computing Machinery, New York, NY, USA, Article 25, 2 pages. https://doi.org/10.1145/3610542.3626124
@@
- Yang et al. (2021) Zhijin Yang, Pengfei Xu, Hongbo Fu, and Hui Huang. 2021. WireRoom: model-guided explorative design of abstract wire art. ACM Trans. Graph. 40, 4, Article 128 (jul 2021), 13 pages. https://doi.org/10.1145/3450626.3459796

### v2

@@ 9.1. Results
The mean SUS score across all participants was 77.3. All participants completed Task 1 successfully within the time limit.
In mechanical testing, all connectors produced in Task 1 reliably supported the bottle. As shown in Fig. 18, participants generated varied designs that successfully conformed to the target bottle.In Task 2, participants created diverse freeform wire structures.
Variation in Connector Arrangements
In Task 1, participants explored different ways of arranging connectors and demonstrated notable variation in how they positioned and combined them. P5 stacked the Bottle Connector above the Table Edge Connector, forming a Z-shaped layered structure (as shown in the Fig. 19 a), while P10 rotated the Bottle Connector so its opening faced inward (as shown in the Fig. 19 b). These variations highlight creative reinterpretations of constrained parts, with participants varying orientation, height, and placement.
Adaptability and Adjustment
Participants also leveraged the adaptability of wire structures. For example, P9 frequently compared their connector to the bottle (as shown in Fig. 20 a) but still produced an oversized design, likely reflecting differences in spatial perception. Nevertheless, P9 successfully adjusted the connector post-fabrication to securely hold the bottle(as shown in Fig. 20 b), illustrating the forgiving nature of wire bending and its tolerance for dimensional inaccuracies (Fig. 20).
Environment-Aware and Expressive Design Strategies
In Task 2, participants displayed greater creative freedom, employing distinct strategies that responded to both environmental context and personal expression. Several designs explicitly incorporated surrounding features: P1 created a hook attached to the blackboard tray ledge (Fig. 21 a), while P8 constructed a structure integrated with a chair surface (Fig. 21 b). These examples illustrate how participants adapted their designs to real-world geometry, leveraging the XR environment’s spatial context. Interestingly, participants exhibited different spatial engagement strategies: some grounded their designs in the physical environment, using surrounding surfaces and contexts to guide their work, while others approached the space more abstractly, positioning forms freely without reference to environmental constraints.
Other participants pursued expressive or artistic goals. P12, an architect, crafted an elegant wire crown using careful symmetry and looping (Fig. 21 c), while P6 designed a flower-like structure—an ambitious concept that would require additional dedicated fixtures to fabricate (Fig. 21 d).
Participants like P1 and P10 transferred structural ideas from Task 1, applying learned connector patterns to their freehand designs. Some struggled with precision due to gesture resolution limitations, which led to larger-than-intended structures. Together, these strategies reveal Y-AR’s potential to support both environment-aware adaptation and creative exploration, while highlighting opportunities to improve fine-detail fabrication support.
9.2. Qualitative Findings
We found four key topics in participants reflection on their experience. The interaction itself, the spatial context brought by XR, the design freedom, and technical limitations.
Intuitive interaction.
Participants described Y-AR as intuitive and easy to learn, with smooth gesture-function mappings that reduced the learning barrier. We computed a learnability subscale score following Lewis and Sauro (2009), who proposed interpreting SUS items 4 and 10 as reflecting learnability. The mean learnability score was 76.0, suggesting participants generally found the system easy to learn.
P4 characterized the core interaction pattern as ”pretty natural” and noted that ”the logic made a lot of sense.” P10 similarly commented that although it usually takes time to understand how to use a new tool, Y-AR was ”intuitive.” P2 emphasized that gesture interactions felt smooth and well-mapped: ”Each gesture has each function, and I can use the bottle to connect or use pen to draw—it’s pretty easy.”
Spatial context and dimensioning.
Several participants highlighted how Y-AR made it easier to design with real-world scale and dimensions in mind. P3 appreciated how the tool automatically aligned and measured distances: ”I can hold on to any object and it kind of maps the distance, which I don’t have to measure manually—that is good.” P10 noted that ”if I’m doing these gestures, it’s very exact… like the dimensions.” In contrast to traditional CAD software, P11 described the difficulty of accurately imagining and measuring objects: ”It’s hard to measure this one and build it precisely, and hard to imagine the real object in [red. traditional 2D CAD] modeling software.”
Expressive freehand design supporting creativity.
Participants valued Y-AR’s ability to support expressive, manual sketching without the constraints of traditional CAD interfaces. P12 described this freedom as a key benefit: ”You can draw anything manually, and you can do what you want to express… you don’t have to be limited by the mouse or by software tools like the…the rules and the circle.” P2 similarly noted that compared to software like Rhino, Y-AR was more convenient for quickly translating ideas into designs: ”I can just think what I want to draw… so it’s very convenient.” Some participants found this approach inspirational; for instance, P7 did not have a clear idea initially but was influenced by the physical environment to design a flower vase: ”Once I had the headset on… I just saw the table was empty… that’s why I built a flower vase with different flowers.”
Limitations.
Despite the overall intuitiveness of Y-AR’s spatial interactions, some participants experienced challenges during initial use, particularly around hand-tracking precision and manipulation accuracy. P4, unfamiliar with XR, described a ”learning curve” and attributed it partly to tracking precision. P5 similarly reported early difficulty with grabbing specific parts of objects but noted that it improved with practice: ”for a specific end, it’s difficult…in the beginning… But then I got used to it.” P6 also described initial interactions as ”shaky,” becoming ”instinctive and natural” over time. These comments highlight that while Y-AR lowers barriers for gesture-based design, new users still require adjustment time to achieve fine control.
9.3. Applications.
We specifically primed participants by asking how and when they would be inclined to use Y-AR in their daily lives. Two key application areas participants surfaced were augmenting existing objects, and creative ideation.
Customization.
Participants widely envisioned Y-AR supporting highly personalized designs tailored to individual needs. P4 described frustrations with off-the-shelf products like phone stands and emphasized the value of being able to ”build something and choose the angle myself.” P6 noted that engineers often require specific components that are unavailable or involve long delivery times, suggesting that Y-AR could help create ”very specific pieces” quickly in maker lab contexts. P8 proposed scenarios such as deployable camping structures or personalized jewelry, where bespoke designs are desirable.
Support ideation.
Several participants identified Y-AR’s potential at the early stages of design work, where quick iteration and open-ended form exploration are essential. Drawing on architectural experience, P3 noted its utility for tasks like steel reinforcement prototyping, describing it as well-suited to ”the design iteration and ideation stage.” P7 and P9 similarly saw value for designers who need to quickly create ”first drafts” or rapidly sketch and visualize ideas before refining them further in traditional CAD tools.
9.4. Feature Requests
Participants provided specific feedback highlighting areas for Y-AR improvement.
Precision and control.
Many participants requested tools to increase the precision of drawing and manipulation. They suggested features such as straight-line drawing similar to ”Microsoft Paint” (P4), linking and constraining lines (P9), and basic shape primitives like circles and rectangles for easy placement (P12). Additional requests included undo functionality (P11), improved object selection in cluttered scenes (P12), and background image tracing to guide accurate sketches (P2). Collectively, these suggestions highlight participants’ desire to complement freeform interaction with accuracy-enhancing mechanisms.
Bridging freeform and modular components.
Several participants expressed interest in seamlessly combining expressive sketches with pre-designed parts, envisioning workflows that mix flexibility and structure. P1 proposed ”connectors that are able to fit your custom drawings,” while others sought CAD-like editing operations such as copy, paste, and delete (P6), and grouping behaviors to keep connected elements together (P7, P12). These ideas reflect participants’ aspirations for workflows that would enable smooth integration between freehand creativity and precise modular assembly.
Expanding ready-made libraries.
Participants noted the usefulness of pre-designed components and requested a broader library of ready-made parts to speed up design. Their suggestions ranged from phone stands and laptop holders to keychains and pin holders (P2, P7, P10, P11, P12). Such libraries would provide convenient starting points for quickly building functional prototypes without requiring all geometry to be drawn from scratch.
Imagining future interaction possibilities.
Some participants offered forward-looking suggestions that would extend Y-AR’s interaction model. P3 envisioned synchronous multi-user collaboration, describing it as ”like Sketchbook, but in a three-dimensional space.” P6 suggested voice commands for generating common primitives, while P12 imagined AI-assisted guidance to help users select appropriate shapes and curves based on their intent. These ideas point to opportunities for expanding Y-AR beyond single-user, gesture-based workflows.
10. Discussion
In this section, we reflect on key insights from the Y-AR project, highlighting its broader implications for XR-CAD tools, wire bending, and fabrication-aware design.
10.1. Fabrication-Aware Design CAD tool for XR
Current CAD/CAM tools for wire bending typically provide low-level machine controls or conversion-based workflows that do not fully leverage the unique material properties and fabrication characteristics of wire structures. These tools often require technical expertise and offer limited design-focused interaction.
Our work explores how fabrication-aware design principles can be integrated into XR-CAD tools for wire bending. Rather than viewing XR tracking inaccuracy as a fundamental flaw, Y-AR demonstrates how fabrication-aware strategies can accommodate these limitations. Specifically, Y-AR leverages material properties such as the springiness of wire to compensate for imprecise input, allowing rough, approximate designs to still result in manufacturable, functional artifacts.
This suggests that XR CAD tools need not replicate the precision-oriented paradigms of traditional CAD systems; instead, they might offer value by supporting design approaches that are inherently robust to variation and uncertainty. Importantly, this approach is extensible beyond wire bending: many CAD-to-fabrication workflows face uncertainties that lead to discrepancies between design and physical outcome. Adaptive structures—those that tolerate or even leverage imprecision—can help bridge this gap by turning fabrication uncertainty into certainty. Future XR-CAD systems could embed tolerance-aware design suggestions or guide users toward structures that naturally accommodate fabrication imprecision, making such adaptability a core design strategy.
10.2. Gesture Discoverability
Another contribution of Y-AR is its support for gesture discoverability. Users did not need to memorize abstract symbolic gestures but could enact natural, mimetic gestures that closely mirror real-world actions (e.g., tracing around a table edge or pointing to a bottle rim).
We observed that participants with little prior XR experience were nonetheless able to quickly adapt to Y-AR’s gesture-based interface. This suggests that gesture interaction provides a natural and accessible way for users to express object relationships in XR environments, similar to metaphors in games (e.g., the intuitive slicing gesture in Fruit Ninja). Our findings indicate that mimetic gesture design allows effective operation without extensive training or memorization.
10.3. Gestural Adaptation in Connector Design
Prior connector design tools, such as AutoConnect (Koyama et al., 2015), primarily rely on predefined geometric shapes (e.g., cylinders, boxes) to generate connection solutions. While this approach appears general, real-world objects exhibit far greater complexity and diversity than these primitives. Most everyday objects feature irregular contours, compound surfaces, or non-standard geometries, limiting the applicability of methods based on standard shapes. Under this shape-to-connector mapping paradigm, the design space becomes constrained by geometric limitations.
Y-AR transforms this landscape through gesture-based interaction. Human hands can naturally trace and conform to complex object contours, from curved bottles to composite surfaces. This shifts the process from ”shape matching” to ”functional exploration.” Moreover, gesture interaction no longer produces singular, deterministic connection solutions. Instead, users can explore multiple connection possibilities for the same object via different gestural paths, gripping strategies, and tracing approaches. For example, users can employ overhead suspension gestures for load-bearing connectors, bottom-up supporting gestures for structural supports, or lateral fixation gestures for clamping mechanisms.
This transition—from constraint to exploration, determinism to diversity—dramatically enriches the design space. Unlike tools enforcing rigid geometric correspondence, Y-AR’s gesture-based approach enables a one-to-many relationship between objects and solutions, where the same physical object yields fundamentally different connector designs depending on the intended interaction mode—each discovered through natural, embodied exploration rather than predetermined geometric analysis.
10.4. Future Work
We found that wire bending is rarely used in isolation but often serves as one component in multi-material fabrication. Wire structures frequently act as frames or supports, combined with other techniques such as 3D printing for joints, laser cutting for panels, or fabric wrapping for surfaces to achieve complete, functional artifacts.
This highlights an opportunity for future CAD tools to support multi-workflow integration, providing an interactive environment where users can holistically plan, visualize, and refine hybrid fabrication workflows. By supporting combinations of additive, subtractive, and formative manufacturing processes, such tools could improve the functionality of their designs.
At the same time, we note that wire bending itself remains underexplored in personal fabrication. Current CAD/CAM tools provide little support for wire-based structures, limiting novices’ ability to explore this design space.
Y-AR represents a first step toward addressing this gap. Future research could expand the wire structure design space by enabling more complex geometries, for example through generative design features that provide sophisticated starting points or automated path optimization tools that reduce manual trial-and-error when planning manufacturable geometries.
11. CONCLUSION
@@
We presented Y-AR, a fabrication-aware design CAD tool for 3D wire bending. This fabrication-aware environment allows designers to create well-designed wire-bent structures. We consider our guidelines for designing for wire bending to be valuable as contributions, independent of the specific implementation of the Y-AR environment. A central insight of our approach is that the springiness of the material can be used to overcome modeling inaccuracies. This allows Y-AR to rely on much less precise input techniques such as mid-air gestures in XR. The XR environment, in turn, supports more context-aware design by helping users create objects that fit their physical surroundings. We believe this core idea will allow developers to build fabrication-aware CAD tools for fabrication techniques that extend beyond wire bending. Our user evaluation indicates that Y-AR allows users to create functional structures with wire bending that would have been hard to accomplish with existing CAM tools.
@@
Y-AR will be available as an open-source tool for other researchers to extend. Future work could explore the integration of novel tool types or the expansion of design capabilities to accommodate fabrication methods beyond wire bending. Related work has shown that wire bending is well equipped to provide structure for other fabrication processes. We are also interested in running more longer-term deployment to see how users benefit from Y-AR over an extended amount of time. And as stated before, the framework of Y-AR, combining XR and springiness, offers a promising foundation for other future fabrication-aware design environments.
@@ Acknowledgements.
We thank the XR Collaboratory at Cornell Tech for their generous sponsorship of this research in the form of the inaugural XR Prototyping Grant. We also acknowledge support from the BURE program at Cornell University. We further thank Pensa Labs for generously providing access to their wire bender, which was essential for prototyping and testing the wire structures presented in this work.References
@@
- Brooke (1996) john Brooke. 1996. SUS: A ’Quick and Dirty’ Usability Scale. In Usability Evaluation In Industry. CRC Press. Num Pages: 6.
@@
- Hsiao et al. (2018) Kai-Wen Hsiao, Jia-Bin Huang, and Hung-Kuo Chu. 2018. Multi-view wire art. ACM Transactions on Graphics 37, 6 (Dec. 2018), 1–11. https://doi.org/10.1145/3272127.3275070
- Huang et al. (2023) Huiguang Huang, Dong-Yi Wu, Hanh Le, Po-Chih Chen, Shih-Syun Lin, and Tong-Yee Lee. 2023. Learning to Generate Wire Sculpture Art from 3D Models. In SIGGRAPH Asia 2023 Posters. ACM, Sydney NSW Australia, 1–2. https://doi.org/10.1145/3610542.3626124
@@
- Koyama et al. (2015) Yuki Koyama, Shinjiro Sueda, Emma Steinhardt, Takeo Igarashi, Ariel Shamir, and Wojciech Matusik. 2015. AutoConnect: computational design of 3D-printable connectors. ACM Trans. Graph. 34, 6, Article 231 (nov 2015), 11 pages. https://doi.org/10.1145/2816795.2818060
- Lewis and Sauro (2009) James R. Lewis and Jeff Sauro. 2009. The Factor Structure of the System Usability Scale. In Human Centered Design, Masaaki Kurosu (Ed.). Springer, Berlin, Heidelberg, 94–103. https://doi.org/10.1007/978-3-642-02806-9_12
@@
- Yang et al. (2021) Zhijin Yang, Pengfei Xu, Hongbo Fu, and Hui Huang. 2021. WireRoom: model-guided explorative design of abstract wire art. ACM Transactions on Graphics 40, 4 (Aug. 2021), 1–13. https://doi.org/10.1145/3450626.3459796
@@
Appendix A Participant Demographics
| ID | Age Group | XR usage | CAD usage | Background |
| P1 | Under 18 | Intermediate | Regularly | FIRST robotics Team |
| P2 | 25–34 | Occasionally | Regularly | Engineer student |
| P3 | 25–34 | Rarely | Regularly | Architecture |
| P4 | 18–24 | Rarely | Never | CS student |
| P5 | 18–24 | Once or twice | Occasionally | CS student |
| P6 | 18–24 | Rarely | Occasionally | CS student |
| P7 | 18–24 | Never | Once or twice | CS student |
| P8 | 25–34 | Rarely | Regularly | Design student |
| P9 | 18–24 | Rarely | Occasionally | Engineer student |
| P10 | 18–24 | Rarely | Once or twice | CS student |
| P11 | 18–24 | Rarely | Never | Design student |
| P12 | 18–24 | Rarely | Regularly | Architecture |